OpenStudy (anonymous):

Put the equation in the form ax^2+bx+c=0 x^2+3x+7=6x+18

OpenStudy (anonymous):

x^2-3x-11=0 a=1 b=-3 c=-11

OpenStudy (anonymous):

Subtract 6x from both sides: x^2 -3x + 7 = 18 Then Subtract 18 from both sides: x^2 - 3x - 11 = 0